Status Report

ISS On-Orbit Status 24 Aug 2002

By SpaceRef Editor
August 24, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

Crew day again began last night at 4:00 pm EDT, ending this morning at 7:30 am.

With a light-duty schedule to complete, ISS residents continued preparing for the spacewalk on Monday morning.  Today’s focus was on Orlan suit fit checks, suited exercise, and communications checks. [After an EVA timeline update with a ground specialist on S-band, CDR Korzun initiated charging for the Russian Glisser-M video camera that is to be used to cover operations at the Japanese MPAC/SEED payloads during the EVA.  He then  set up the CT (communications and tracking) configuration for EVA, and, with FE-2 Sergey Treschev, checked out the Orlan systems functions.  Wearing their biomedical instrumentation belts, the cosmonauts then tested biomedical comm for vital signs monitoring.  After donning equipment, they ingressed the suits through their open backs, then closed the backpacks.  During the following suited exercise, the Orlans and their BSS controls were checked out again, tested for leaks and subjected to fit checks at reduced suit pressure.  Valery and Sergey then tried out their ability to move about in the Orlans and evaluated their spatial orientation and positioning in the DC-1 while testing BITS telemetry connections for the EVA.  At day’s end, the CDR terminated Glisser battery charging.]

FE-1 Peggy Whitson refreshed her familiarity with the DOUG (dynamic operational ubiquitous graphics) application, a sophisticated software program on a standalone PC at the robotics workstation.  It provides a birdseye-view graphical image of the SSRMS arm, showing its real-time location and configuration on the display during its operation.

Later, she terminated the recharging of the EMU batteries started yesterday and completed the daily routine maintenance tasks of SOSh life support systems servicing, Lab payloads status checkup, and IMS delta file preparation.

All crewmembers performed their daily physical exercise program on TVIS, CEVIS, VELO with load trainer, and RED.

The crew was thanked for their downlinked TVIS video of yesterday (8/23).  MCC-H specialists are requesting more video, preferably covering a full exercise period, since this would be very helpful for Engineering to understand exactly how TVIS is utilized. [Specifically, the video should show the exercise portion for at least 15 minutes duration, including a minimum of five minutes of harness donning, doffing, and use of the control panel.  The picture frame should include the entire body, from head to feet, and camera angle should be from the side as much as possible since this provides maximum information about running style and how the foot strikes on the belt.]

A PFC (private family conference) via Russian ground sites for FE-2 Treschev had to be cancelled when the ground party at MCC-M could not be contacted.  FE-1 Whitson had her weekly PFC via S-band.

Today’s targets for the CEO program were Baghdad, Iraq (nadir view and slightly left of nadir.  Crew was to try for margin of built-up areas. ESC [electronic still camera] was requested), Eastern Mediterranean Aerosols (Aegean Sea [on left] receives smog from both northern Italy and eastern Europe [via the Danube valley and Black Sea]. Handheld photos reveal smog flow trajectory well), Western Mediterranean Aerosols (crew was to continue to shoot oblique views for smog loadings.  The Med basin concentrates smog whenever winds blow south), Algiers, Algeria (nadir view; ESC requested), Rome, Italy (nadir view; look inland from the coast.  ESC requested), Central Europe floods (excellent crop of recent detailed ESC images of this flood event were received on the ground.  The crew was asked to document the declining phase: looking nadir for larger rivers in the Czech Republic and southern Poland), Amazon fires (numerous fires over a remarkably clear Amazon basin.  Low sun angles assist documentation of variable smoke loadings in the atmosphere), Taiwan Smog (weeks-long cloud cover has moved away.  Crew to photograph possible smog buildup in strait between island and mainland China), and Black Hills fires (eleven-thousand acres blackened after possible arson, four miles north of Mt Rushmore [Black Hills visible as an oval darker region in the plains well east of the Rocky Mt front].  Crew to look right for smoke and new dark fire scar).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 6:05 am EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (32-amp mode), on backup pump.  Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5, i.e., 10-min. cycle time (vacuum pump failed).  U.S. CDRA CO2 scrubber is on Override.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
  • SM Working Compartment:  Pressure (mmHg) — 760, temperature (deg C) — 26.9, ppO2 (mmHg) — 157.2, ppCO2 (mmHg) — 2.6.
  • SM Transfer Compartment:  Pressure (mmHg) — 751, temperature (deg C) — 19.8.
  • FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C) — 20.0.
  • Node:  Pressure (mmHg) — 752.79, temperature (deg C) — 23.0 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 754.92, temperature (deg C) — 25.0, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 754.92, temperature (deg C) — 26.2; shell heater temp (deg C) — 26.6, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 21.9
  • PMA-2:  Shell heater temp (deg C) — 19.0

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas assembly) were replaced, but some more work needs to be done).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and BGA 4B in Autotrack mode (solar-tracking).
  • SM batteries:   Battery #1 is off-line; all other batteries (7) are in “Full Charge” mode.
  • FGB batteries:  Battery #5 is off-line; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:

  • C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is operational; PL-2 MDM is Off.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line.
  • State vector — US GPS (SIGI string 1)
  • Attitude — Russian segment
  • Angular rates — US RGA1 (rate gyro assembly 1)

Communications & Tracking Systems:

  • All Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data grapple fixture 1), in EVA-7 viewing position, with Keep Alive power on both strings (based on MBS).
  • MBS: Keep Alive power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 6:25am EDT [= epoch]):

  • Mean altitude — 392.2 km
  • Apogee — 403.5 km
  • Perigee — 380.8 km
  • Period — 92.4 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0016764
  • Orbits per 24-hr. day — 15.58
  • Altitude decrease — 250 m (mean) in last 24 hours
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 21467
  • Current Flight Attitude — LVLH (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: 7.8 deg, roll: 0 deg]).  Next change will be back to XPOP on 8/26.

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.